OpenStudy (anonymous):

Maximum Value of |Z| if |Z+4/Z | = 3

OpenStudy (anonymous):

i used this relation |Z1+Z2| < = |Z1| + |Z2| ...But my Result is wrong..:_

OpenStudy (shubhamsrg):

your attempt seems fine. can you show your work ?

OpenStudy (anonymous):

Finally after Solving i Got an Equation like this |Z|^2 - 3|Z| + 4 > = 0

OpenStudy (shubhamsrg):

these seems to be always > 0 ?

OpenStudy (shubhamsrg):

it should have a minimum value then unless some mistake ?

OpenStudy (anonymous):

Sorry.....Did nt get u....

OpenStudy (anonymous):

i think u mean it shuld be < = 0 for maximum ryt ?

OpenStudy (shubhamsrg):

the given quadratic eqn can not be simplified. it must be a parabola with mouth open upwards such parabolas have minimum vale

OpenStudy (shubhamsrg):

value*

OpenStudy (anonymous):

So..Hw can We approach this ?

OpenStudy (shubhamsrg):

|Z+4/Z | <= |z| + 4/|z| 3 <= |z| + 4/|z| |z|^2 - 3|z| + 4 >=0 hmm this is always greater than 0 for any value of |z| you should be asking minimum value of |z| ?

OpenStudy (anonymous):

there are no real values for which \[\left|z+{4\over z}\right|=3\]

OpenStudy (anonymous):

\[ |z|^2+8+{16\over|z|^2}=9\\ |z|^4-|z|^2+16=0 \]

OpenStudy (anonymous):

the solution for this system says that the "z" is a complex number